Given the values of ΔHorxn, ΔSorxn, and T below, determine ΔSuniv.

     ΔHorxn= -95 kJ , ΔSorxn= -157 J/K , T= 855 K

(HINT: Use the equation on page 876,  LaTeX: \DeltaΔSuniv  = LaTeX: \DeltaΔS -  LaTeX: \DeltaΔH / T ).  

Group of answer choices

-45.9 J/K

-111.1 J/K

-157 J/K

+238.7 J/K

-268.11 J/K
Answer #1

Given:

ΔH = -95 KJ

= -9.5*10^4 J

T = 855 K

use:

ΔS Surr = -ΔH/T

= -(-9.5*10^4)/855

= 1.111*10^2 J/K

ΔSuniv = ΔS Sys + ΔS Surr

= -157 J/K + 1.111*10^2 J/K

= -45.89 J/K

Answer: -45.9 J/K